OpenStudy (anonymous):

Find the area of a rectangular object which has a length of 2√3 and a width of 3√6. Explain, in complete sentences, how you arrived at the simplified answer and give the final solution in simplified radical form. What type of object might this be?

OpenStudy (harsimran_hs4):

what is the area of a rectangular object ?

OpenStudy (anonymous):

b x h

OpenStudy (harsimran_hs4):

l * b (l = length b = breadth) so putting the values given in the ques. what expression do you get ?

OpenStudy (anonymous):

2√3 x 3√6 ?

OpenStudy (harsimran_hs4):

yes, so \[3\sqrt{6} = 3 \sqrt{2\times3} = 3 \sqrt{2}\sqrt{3}\]

OpenStudy (harsimran_hs4):

so 2sqrt(3) * 3sqrt(2)*sqrt(3) = 2*3*(sqrt(3) * sqrt(3))*sqrt(2)

OpenStudy (harsimran_hs4):

clear?

OpenStudy (anonymous):

what do you do with the 2√3 ?

OpenStudy (harsimran_hs4):

2sqrt(3) = 2* sqrt(3) so all i have done is paired this sqrt(3) with other one

OpenStudy (anonymous):

are the * = multiplication ?

OpenStudy (harsimran_hs4):

yes

OpenStudy (harsimran_hs4):

generally +/ - = addition/subtration * = multiplication / = divide ^ = raised to power

OpenStudy (anonymous):

ooh so 2√3 x 3√2 x √3 = 2 x 3 x √3 x √3 x √2 is the answer ?

OpenStudy (harsimran_hs4):

no we can simplify more sqrt(3) * sqrt(3) = 3 and 2*3 = 6 so 6 * 3 * sqrt(2) 18*sqrt(2) \[18\sqrt{2}\] is the answer

OpenStudy (anonymous):

okay thank you !